# Weinstein exactness of nearby Lagrangians and the Lagrangian $C^{0}$ flux conjecture

### v2

@@ Abstract
This paper is concerned with the following problem: given a Lagrangian $L$ and a Hamiltonian diffeomorphism $\varphi$ such that $\varphi(L)$ is in a small neighbourhood $U$ of $L$, does there exist another Hamiltonian isotopy from $L$ to $\varphi(L)$ supported in $U$? On one side, we construct an irrational counterexample in any symplectic manifold of dimension at least six. On the other side, we answer an a priori weaker form of the question by the positive in many cases when $L$ satisfies some rationality condition. The techniques that we develop also have applications for the Lagrangian counterpart of the $C^{0}$ flux conjecture. In turn, these results have many applications, in particular, to understand $C^{0}$-rigidity phenomena of Hamiltonian diffeomorphisms and the space of Lagrangians with a given rationality constant.
@@ 1 Introduction
Note: For the sake of conciseness, we will refer throughout the paper to a closed connected Lagrangian submanifold of a connected symplectic manifold without boundary as a “Lagrangian in a symplectic manifold.”
@@
To do this, we first fix a metric $g$ on the underlying manifold $L$. Recall that the Weinstein neighbourhood theorem ensures that there exist $r>0$ and a symplectomorphism $\Psi:D^{*}_{r}L\to\mathcal{W}_{r}(L)$ from the codisk bundle of $L$ of radius $r$ to a neighbourhood $\mathcal{W}_{r}(L)$ of $L$ in $M$ which maps the 0-section to $L$.
Therefore, understanding $\mathcal{L}{\mathrm{Ham}}(L)$ locally is intimately related to the nearby Lagrangian conjecture (or NLC for short), which completely characterizes Lagrangians which are in the Hamiltonian orbit of the 0-section in $T^{*}L$. Indeed, it states that those are precisely the exact Lagrangians. It is known to hold for $S^{1}$, $S^{2}$ [Hin04], $\mathds{R}\mathrm{P}^{2}$ [HPW16, Ada22], and $\mathds{T}^{2}$ [RGI16]. Without restriction on the diffeomorphism type, the most advanced result in the direction of the NLC states that the natural projection $\pi:T^{*}L\to L$ induces a (simple) homotopy equivalence between any exact closed Lagrangian and the 0-section [AK18]. This latter result will play a crucial role in our study of the local structure of $\mathcal{L}{\mathrm{Ham}}(L)$
Inspired by this conjecture, we propose that if $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L)$ is close to $L$, then there is an accordingly small Hamiltonian isotopy from $L$ to $L^{\prime}$. More precisely, we make the following conjecture.
Conjecture A (Strong conjecture)
@@
That this conjecture holds would imply that the Hamiltonian orbit $\mathcal{L}{\mathrm{Ham}}(L)$ of a Lagrangian $L$ is locally path connected via Hamiltonian isotopies. Although this statement is new in the Lagrangian context, there are some results towards its Hamiltonian counterpart. More precisely, the group $\mathrm{Ham}_{c}(M)$ of compactly supported Hamiltonian diffeomorphisms of a symplectic manifold $M$ is locally path connected in the $C^{0}$ topology (via Hamiltonian isotopies) if $M$ is a closed surface or the open ball $B^{2n}$. The former case follows from Fathi’s work on homeomorphisms preserving a volume form [Fat80] and the folkloric fact that a path of such homeomorphisms on a closed surface can be $C^{0}$ approximated by a path of symplectomorphisms (see [Oh06] for a proof). The latter case was proved by Seyfaddini [Sey13].
Note that the local path connectedness of $\mathrm{Ham}(M)$ implies Conjecture A for graphs of symplectomorphisms of $M$. However, even if Conjecture A holds for all graphs in $M\times M$, it does not imply local path connectedness of $\mathrm{Ham}(M)$, since the Hamiltonian isotopy given by the conjecture is not necessarily through graphs.
In this work, we prove the existence of neighbourhoods of local exactness for several classes of Lagrangians, by which we mean a Weinstein neighbourhood $\mathcal{W}(L)$ of a Lagrangian $L$ so that any Lagrangian Hamiltonian isotopic to $L$ included in this neighbourhood is exact in $\mathcal{W}(L)$. When the NLC is known to hold for $L$, we can then deduce that $L$ satisfies the strong conjecture above.
This indicates that it is reasonable to believe that the strong conjecture is extremely hard to prove in general. However, the following weaker form of the conjecture holds and might be more easily provable in full generality.
Conjecture B (Weak conjecture)
@@
The Lagrangians which admit a neighbourhood of local exactness as described above obviously satisfy the weak conjecture. In fact, such Lagrangians must intersect in at least $\sum_{i=0}^{n}\beta_{i}(L)\geq 2$ points, where $\beta_{i}(L)$ is the $i$-th Betti number of $L$.
Actually, our methods allow us to prove that a large class of Lagrangians satisfy a slightly strengthened version of the weak conjecture, namely that if $L^{\prime}$ is the image of $L$ under any symplectomorphism and $L^{\prime}\subseteq U$, then $L\cap L^{\prime}\neq\varnothing$. Note that we do not conjecture that this always holds (since it obviously does not). In what follows, we will refer to satisfying the weak conjecture (respectively its symplectic version) as having a neighbourhood of Hamiltonian nondisplacement (respectively of symplectic nondisplacement).
@@ 1.1 Main results
Our first result provides a counterexample to Conjectures A and B (and to Conjecture C below, as well as its $C^{1}$ variant). This shows that there is no hope to prove the above conjectures in full generality.
@@
    is not closed in Hausdorff topology inside the set of Lagrangian tori,
@@
    admits arbitrarily Hausdorff-close disjoint elements.
@@
We make the crucial observation that these Lagrangian tori are not rational, and we turn to specific families of Lagrangians which do satisfy either one or both of the conjectures above.
@@ 1.1.1 Existence of nondisplacement neighbourhoods
We start with a simple exercise, which inspires our approach to the above conjectures.
Proposition 2
Conjecture A holds for any closed embedded curve $L$ in an orientable surface $M$.
Its proof, detailed in Appendix A, can be roughly summarized as follows.
Sketch of proof.
Fix a Weinstein neighbourhood $\Psi:D^{*}_{r}S^{1}\to\mathcal{W}_{r}(L)$ of $L$, and let $L^{\prime}=\varphi(L)\subseteq\mathcal{W}_{r}(L)$ for some symplectomorphism $\varphi$ of $M$. Define
where we set $\tau=+\infty$ if $\omega(H_{2}(M,L;\mathds{Z}))=0$.
If $\tau>0$, the proof then follows in two steps.
  1. (1)
    Show that, if $r>0$ is such that $\mathcal{W}_{r}(L)$ has area smaller than $\tau$, then $L^{\prime}$ must be weakly exact in $\mathcal{W}_{r}(L)$.
  2. (2)
    Show that if $L^{\prime}$ is weakly exact in $\mathcal{W}_{r}(L)$, then it must in fact be exact.
Since the nearby Lagrangian conjecture is known to hold for the circle, this yields Conjecture A when $\tau>0$.
If $\tau=0$, simple geometric considerations due to the low dimension of the situation allow us to easily adapt the proof above. ∎
Inspired by this simple case, we develop a criterion which should ensure the existence of neighbourhoods of nondisplacement for Lagrangians in arbitrary dimensions. Namely, we define a notion of homological rationality, or $H$-rationality for short.
Definition
Let $L$ be a Lagrangian of a symplectic manifold $(M,\omega)$. We say that $L$ is $H$-rational in $(M,\omega)$ if $\omega\big{(}H_{2}(M,L;\mathds{Z})\big{)}=\tau\mathds{Z}$ for some $\tau\geq 0$. We then call $\tau$ the $H$-rationality constant of $L$. When $\tau=0$, we say that $L$ is $H$-exact in $(M,\omega)$.
Note that $H$-rationality (respectively $H$-exactness) is the homological version of the usual notion of rationality (respectively weak exactness). In fact, as we shall see in Section 3.2 below, those conditions are equivalent in many important cases, e.g. when $\pi_{1}(M)=0$.
@@
We now adapt the two steps of the 1-dimensional case to higher dimensions. First, we prove the existence of neighbourhoods of homological exactness for several classes of Lagrangians — this is Theorem 3 below. Second, we show that such neighbourhoods lead to Weinstein neighbourhoods of exactness for Hamiltonian isotopic Lagrangians. We also get such neighbourhoods for Lagrangians with a given $H$-rationality constant under an extra homological condition — this is the content of Theorem 4.
@@
Suppose that $L$ is a Lagrangian submanifold of $(M,\omega)$ which satisfies one of the following:
  1. (a)
    the underlying manifold has $H_{1}(L;\mathds{R})=\mathds{R}$ and admits a Lagrangian embedding in a Liouville domain $W$ with $SH(W)=0$,
  2. (b)
    $L$ is a Klein bottle.
For every $\tau\geq 0$, there exists a Weinstein neighbourhood $\mathcal{W}(L)$ of $L$ such that all $L^{\prime}\in\mathcal{L}(L,\tau)$ included in $\mathcal{W}(L)$ is $H$-exact in $\mathcal{W}(L)$.
@@
Let $L$ be a $H$-rational Lagrangian submanifold of $(M,\omega)$ and let $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L)$ be a Lagrangian included in a Weinstein neighbourhood $\mathcal{W}_{r}(L)$ of size $r>0$ such that $L^{\prime}$ is $H$-exact in $\mathcal{W}_{r}(L)$. Then, for a maybe smaller $r$, $L^{\prime}$ is exact in $\mathcal{W}_{r}(L)$.
@@
Remarks 1.
In many instances, the constraints on $L$ and $L^{\prime}$ can actually be weakened. We refer the interested reader to Section 3.2 for specifics.
@@
A $H$-rational Lagrangian $L$ satisfying the assumptions of Theorem 3 admits a Hamiltonian non-displacement neighbourhood. If furthermore $H_{1}(L;\mathds{R})\to H_{1}(M;\mathds{R})$ is zero, then it also admits a neighbourhood of symplectic non-displacement.
In other words, such a Lagrangian $L$ satisfies Conjecture B (or its slightly strengthened symplectic version). Furthermore, if the Nearby Lagrangian Conjecture holds in $T^{*}L$, we also get Conjecture A.
Remarks 2.
Theorem 3 above also holds for $n$-dimensional Lagrangian tori. This can be proven along the same lines as for case (a) by bounding our variant of the Cieliebak–Mohnke capacities by the McDuff–Siegel higher capacities. However, the latter being much harder to handle, the proof becomes much more involved and will appear in further work.
@@
    In dimension 2, we can prove that
    1. (a)
      A displaceable rational Lagrangian 2-torus in a simply-connected Darboux chart admits a neighbourhood of symplectic nondisplacement, i.e. satisfies the symplectic version of Conjecture B.
    2. (b)
      A nondisplaceable Lagrangian 2-torus $L$ whose fundamental class $[L]$ does not vanish in $H_{2}(M;\mathds{Z})$, satisfies Conjecture A.
    Because the proofs of these results do not fit in the general framework developed here, we prove them in Appendix B, where we also state them precisely.
@@
    In their work in progress on the $C^{0}$ flux conjecture for Hamiltonian diffeomorphisms, Atallah and Shelukhin [AS24a] get a similar result of local exactness for graphs in $M\times M$ of $C^{0}$-small Hamiltonian diffeomorphisms for $M$ closed. In this case, they do not require that $M$ be rational, contrary to our setup.
  3. 3.
    As we were finalizing this paper, Atallah and Shelukhin also informed us [AS24b] that they had results similar to Corollary 5. In particular, they established a version of Theorem 4 whose proof is strikingly similar to ours. They also proved a version of Theorem 3 for rational Lagrangians $L$ of the form $\mathds{T}^{n}\times Q$, where $Q$ is simply connected, under the assumption that $\pi_{1}(L)\to\pi_{1}(M)$ is zero. It appears that combining their methods and ours leads to the following fact: if $L$ and $L^{\prime}$ satisfy the conclusion of Theorem 3, then so does $L\times L^{\prime}$. We want to investigate this exciting direction together.

### v3

# Weinstein exactness of nearby Lagrangians: towards the Lagrangian $C^{0}$ flux conjecture

@@ Abstract
We address the following problem: if a Hamiltonian diffeomorphism maps a Lagrangian submanifold $L$ to a small Weinstein neighborhood of $L$, is the image necessarily Hamiltonian isotopic to $L$ inside that neighborhood? On one hand, we show that the question can have a negative answer in any symplectic manifold of dimension at least six. On the other hand, we answer an a priori weaker form of the question in the positive in various cases when $L$ satisfies a rationality condition: we prove that the image of $L$ is often exact inside the Weinstein neighborhood. We provide applications to the Lagrangian counterpart of the $C^{0}$ flux conjecture, to $C^{0}$-rigidity phenomena of Hamiltonian diffeomorphisms, and to topological properties of spaces of Lagrangians with the same rationality constraint. Moreover, we state and prove cases of an analogue of Viterbo’s spectral norm conjecture for non-exact Lagrangians; in the process, we make progress on an old question of Viterbo regarding integer difference vectors between points of Lagrangians.
@@ 1 Introduction
Note that for the sake of conciseness, we will refer throughout the paper to a closed connected Lagrangian submanifold of a connected symplectic manifold without boundary as a “Lagrangian in a symplectic manifold”.
To enable our study, we first fix a metric $g$ on the underlying manifold $L$. Recall that the Weinstein neighbourhood theorem ensures that there exist $r>0$ and a symplectomorphism $\Psi:D^{*}_{r}L\to\mathcal{W}_{r}(L)$ from the codisk bundle of $L$ of radius $r$ to a neighbourhood $\mathcal{W}_{r}(L)$ of $L$ in $M$ which maps the 0-section to $L$.
Therefore, understanding $\mathcal{L}{\mathrm{Ham}}(L)$ locally is intimately related to the nearby Lagrangian conjecture (or NLC for short), which completely characterizes Lagrangians which are in the Hamiltonian orbit of the 0-section in $T^{*}L$. Indeed, it states that those are precisely the exact Lagrangians. It is known to hold for $S^{1}$, $S^{2}$ [Hin04], $\mathds{R}\mathrm{P}^{2}$ [HPW16, Ada22], and $\mathds{T}^{2}$ [RGI16]. Without restriction on the diffeomorphism type, the most advanced result in the direction of the NLC states that the natural projection $\pi:T^{*}L\to L$ induces a simple homotopy equivalence between any exact closed Lagrangian and the 0-section [AK18]. This latter result will play a crucial role in our study of the local structure of $\mathcal{L}{\mathrm{Ham}}(L)$.
Inspired by this conjecture, we propose that if $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L)$ is close to $L$, then there is an accordingly small Hamiltonian isotopy from $L$ to $L^{\prime}$. More precisely, we consider the following speculation.
Speculation A
@@
If this holds for $L$, then its Hamiltonian orbit $\mathcal{L}{\mathrm{Ham}}(L)$ is locally path connected via Hamiltonian isotopies. Although this statement is new in the Lagrangian context, there are some results towards its Hamiltonian counterpart. More precisely, the group $\mathrm{Ham}_{c}(M)$ of compactly supported Hamiltonian diffeomorphisms of a symplectic manifold $M$ is locally path connected in the $C^{0}$ topology (via Hamiltonian isotopies) if $M$ is a closed surface or the open ball $B^{2n}$. The 2-dimensional case case follows from Fathi’s work on homeomorphisms preserving a volume form [Fat80] and the folkloric fact that a path of such homeomorphisms on a closed surface can be $C^{0}$ approximated by a path of symplectomorphisms, see [Sey13] for a proof. The case of the open ball was proved by Seyfaddini, also in [Sey13].
Note that the local path connectedness of $\mathrm{Ham}(M)$ does not yield Speculation A for all Lagrangians of $M$. Theorem 1 below shows that, for instance in the standard symplectic ball $M$ of dimension $6$ or greater, there exist certain irrational tori for which Speculation A fails. It does, however, imply Speculation A for graphs of Hamiltonian symplectomorphisms of $M$. Conversely, note that even if Speculation A holds for all graphs in $M\times M$, it does not imply local path connectedness of $\mathrm{Ham}(M)$, since the Hamiltonian isotopy given by the speculation need not be through graphs.
In the positive direction, we prove the existence of neighbourhoods of local exactness for several classes of Lagrangians, by which we mean a Weinstein neighbourhood $\mathcal{W}(L)$ of a Lagrangian $L$ so that any Lagrangian Hamiltonian isotopic to $L$ contained in this neighbourhood is exact in $\mathcal{W}(L)$. When the NLC is known to hold for $L$, we can then deduce that $L$ satisfies Speculation A. This indicates, however, that Speculation A is likely extremely hard to prove in great generality.
However, the following weaker form of the speculation holds in greater generality and is easier to prove.
Speculation B
@@
The Lagrangians which admit a neighbourhood of local exactness as described above obviously satisfy that weaker speculation. In fact, such Lagrangians must intersect $L$ in many points: for instance, if the intersection is transverse, then there are at least $\sum_{i=0}^{n}\beta_{i}(L)\geq 2$ intersection points, where $\beta_{i}(L)$ is the $i$-th Betti number of $L$.
Actually, our methods allow us to prove that a large class of Lagrangians satisfy a slightly strengthened version of Speculation B, namely that if $L^{\prime}$ is the image of $L$ under any symplectomorphism and $L^{\prime}\subseteq U$, then $L\cap L^{\prime}\neq\varnothing$. In what follows, we will refer to satisfying Speculation B (respectively its symplectic version) as having a neighbourhood of Hamiltonian nondisplacement (respectively of symplectic nondisplacement).
To give an idea of the methods we use to prove our results, we introduce a Viterbo-style conjecture. To do so, we need the following rationality notions.
Definition
Let $L$ be a Lagrangian of a symplectic manifold $(M,\omega)$. We say that $L$ is $H$-rational in $(M,\omega)$ if the group of relative periods $\omega(H_{2}(M,L;\mathds{Z}))=\tau_{L}\mathds{Z}\subset\mathds{R}$ is a discrete subgroup. We call $\tau_{L}\geq 0$ the $H$-rationality constant of $L.$ We say that $L$ is $H$-exact if $\tau_{L}=0$.
Restricting to disk classes, that is, the image $H^{D}_{2}(M,L;\mathds{Z})$ of $\pi_{2}(M,L)$ under the Hurewicz morphism, $L$ is called rational if $\omega(H^{D}_{2}(M,L;\mathds{Z}))\subset\mathds{R}$ is a discrete subgroup — weakly exact if it vanishes — and we define the rationality constant of $L$ analogously.
We shall see in Section 3.2 below that the homological and homotopical conditions are equivalent in many important cases, e.g. when $\pi_{1}(M)=0$. With these definitions in mind, we make the following conjecture, which we prove in several cases below.
Conjecture C
Let $L$ be a closed connected manifold. Suppose that $K$ is an $H$-rational Lagrangian inside $D^{*}_{r}L$ such that the map $(\pi_{K})_{*}:H_{1}(K;\mathds{R})\to H_{1}(L;\mathds{R})$ induced by $\pi_{K}=\pi|_{K}$ is not surjective. Here, $\pi:T^{*}L\to L$ denotes the canonical projection. Then, the $H$-rationality constant $\tau_{K}$ of $K$ satisfies
$\tau_{K}\leq Cr$
for a constant $C$ depending only on $L$ and the choice of an auxiliary metric on it.
Remark 1.
We note that the non-surjectivity of $(\pi_{K})_{*}$ can be replaced by the inequality of first Betti numbers $b_{1}(K;\mathds{R})\leq b_{1}(L;\mathds{R})$.
One can also pose a homotopy version of Conjecture C, where one replace either $H$-rational with rationality or $H_{1}(\,\cdot\,;\mathds{R})$ with $\pi_{1}$ — or both. We will also explore such variant below.
We remark that it would be interesting to see if there is a version of Conjecture C that also applies to irrational Lagrangians, with the role of $\tau_{K}$ being replaced by $\hbar(K,J)=\min\{\omega(u)>0\}$, where the minimum runs over all the non-constant $J$-holomorphic disks $u:\mathbb{D}\to M$ with boundary on $K$ for a fixed compatible almost complex structure $J.$ We leave this question for future work.
@@ 1.1 Main results
The following result provides a counterexample to Speculations A and B (and to Speculation D below, as well as its $C^{1}$ variant). It is not, however, a counterexample to the Viterbo-type conjecture. This shows that there is no hope to prove the above speculations in full generality, and that a rationality condition like above is required.
@@
    admits arbitrarily Hausdorff-close disjoint elements,
@@
    is not closed in Hausdorff topology inside the set of Lagrangian tori.
@@
Again, we make the crucial observation that these Lagrangian tori are not rational, and therefore do not contradict our positive results below.
1.1.1 Rationality bound on nearby Lagrangians
We first describe positive results related to Conjecture C.
Theorem 2
Conjecture C holds for $L=L_{0}\times L_{1}\times\dots\times L_{k}$, where $H_{1}(L_{0};\mathds{R})=0$, and $L_{i}$, $i\geq 1$, satisfies $H_{1}(L_{i};\mathds{R})=\mathds{R}$ and admits a Lagrangian embedding in a Liouville domain $W_{i}$ with $SH(W_{i})=0.$ Furthermore, if $L$ is any manifold covered by a closed connected manifold $L^{\prime}$ for which Conjecture C holds, then it also holds for $L.$
Note that, in the above statement, $L_{0}$ may be a point.
Remark 2.
The homotopy version of Theorem 2 also holds, that is, when $H$-rationality is replaced with rationality on disks and $H_{1}$ with $\pi_{1}$. When $L$ has the diffeomorphism type of $L_{0}\times\mathds{T}^{m}$ with $H_{1}(L_{0};\mathds{R})=0$, it is possible, via an alternative argument, to prove a stronger version of the conjecture where $K$ is only assumed to be rational on disks and $(\pi_{k})_{*}$ is not surjective on homology; this is proven in Section 4.5, as Theorem 39.
Theorem 39 and an examination of its proof implies the following result, which is a partial answer to an old question of Viterbo. Consider the subset $X\subset\mathds{R}^{n}$ given by $X=\{(p_{1},\ldots,p_{n})\;|\;\min_{j}|p_{j}|\leq 1\}.$
Proposition 3
Let $K$ be a rational Lagrangian inside $\mathds{C}^{n}$ with rationality constant $\tau_{K}>2.$ If $K$ is contained in $\mathds{R}^{n}+iX$ then there exist two distinct points on $K$ whose difference vector is real with integer coordinates.
To provide context, Viterbo’s original question consists in proving the similar statement that there exists a constant $A\geq 1$ such that every rational Lagrangian $K$ inside $\mathds{C}^{n}$ with rationality constant $\tau_{K}>A$ must have two distinct points whose difference vector has Gaussian integer coordinates. This is equivalent to finding the maximal rationality constant of a nullhomotopic rational Lagrangian in the standard torus $\mathds{C}^{n}/(\mathds{Z}^{n}+i\mathds{Z}^{n})$. We prove a stronger statement with the difference vector being also real, under the additional assumption that $K$ is contained in $\mathds{R}^{n}+iX$.
@@ 1.1.2 Existence of nondisplacement neighbourhoods
We now study Speculations A and B. First, we prove the existence of neighbourhoods of homological exactness for several classes of Lagrangians by reducing it to Conjecture C — this is Theorem 4 below. Second, we show that such neighbourhoods lead to Weinstein neighbourhoods of exactness for Hamiltonian isotopic Lagrangians. We also get such neighbourhoods for Lagrangians with a given $H$-rationality constant under an extra homological condition — this is the content of Theorem 5.
@@
Suppose that $L$ is a Lagrangian submanifold of $(M,\omega)$ for which Conjecture C holds. Then, for each $\tau\geq 0$, there exists a Weinstein neighbourhood $\mathcal{W}(L)$ of $L$ such that all $L^{\prime}\in\mathcal{L}(L,\tau)$ included in $\mathcal{W}(L)$ is $H$-exact in $\mathcal{W}(L)$.
@@
Let $L$ be a $H$-rational Lagrangian in $(M,\omega)$ and let $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L)$ be a Lagrangian included in a Weinstein neighbourhood $\mathcal{W}_{r}(L)$ of size $r>0$ such that $L^{\prime}$ is $H$-exact in $\mathcal{W}_{r}(L)$. Then, for a maybe smaller $r$, $L^{\prime}$ is exact in $\mathcal{W}_{r}(L)$.
@@
The existence of nondisplacement neighbourhoods follows, as a direct combination of Theorems 4 and 5.
@@
A $H$-rational Lagrangian $L$ in a symplectic manifold $M$ satisfying the assumptions of Theorem 2 admits a Hamiltonian nondisplacement neighbourhood. If furthermore the map $H_{1}(L;\mathds{R})\to H_{1}(M;\mathds{R})$ is zero, then it also admits a neighbourhood of symplectic nondisplacement.
In other words, such a Lagrangian $L$ satisfies Speculation B (or its slightly strengthened symplectic version). Furthermore, if the nearby Lagrangian conjecture holds in $T^{*}L$, we also get Speculation A.
Remark 3.
We expect that one could use the McDuff–Siegel higher capacities to obtain Theorem 4 for certain $K(\pi,1)$ spaces, including the $n$-torus. Moreover, we expect a slightly different approach to produce a class of examples invariant under all maps of non-zero degree, rather than coverings. Finally, we expect to extend our results to a large class of fibrations over tori. However, the arguments are quite involved and will be investigated in further work.
@@
    There exists yet another approach to the above questions for $L=\mathds{T}^{2}$ based on the resolution [RGI16] in this case of the nearby Lagrangian conjecture. Because the proofs of these results do not fit in the general framework developed here, we prove them in Appendix B, where we also state them precisely. Moreover, in Appendix C we present a standalone proof of local exactness for Lagrangian Klein bottles in the cotangent bundle of a Klein bottle.
@@
    In their work in progress on the $C^{0}$ flux conjecture for Hamiltonian diffeomorphisms [AS24] the first and last named authors get a similar result of local exactness for graphs in $M\times M$ of $C^{0}$-small Hamiltonian diffeomorphisms for $M$ closed. In this case, there is no requirement that $M$ be rational, contrary to the setup of the present work.